OpenStudy (anonymous):

In two years, Jack will be twice as old as Jen. The sum of Jack's age and three times Jen's age is 67. How old is Jack?

OpenStudy (amistre64):

jack + 3(jen) = 67 jack = 67 -3(jen). (jack +2) = 2(jen)....